OpenStudy (anonymous):

Is this correct? Convert the equations you arrived at in question 2 and question 4 into slope-intercept form. Question 2: y - 5 = 1.5(x - 2) y = 1.5x - 3 + 5 y = 1.5x + 2 Question 4: y - 8 = 1.5(x - 4) y = 1.5x - 4 + 8 y = 1.5x + 4

OpenStudy (nurali):

Question 4 wrong y - 8 = 1.5(x - 4) y = 1.5x - 6 + 8 y = 1.5x + 2

OpenStudy (anonymous):

@Nurali thank you

OpenStudy (nurali):

My Pleasure.